If you have a square that is 6 ft on each side and you want to make a triangle with the same sq. footage, how would you and what would be the length of each side. ( length of diagonal line)
Answer by wgomero(28)

area of the square = 6ft * 6ft = 36ft^2
since you want the area of the triangle the same as the square then:
Triangle area = (b*h)/2
36ft^2 = (6ft* h)/2
36ft^2 = 3ft*h
(36ft^2)/3 = h
h = 12ft
so the two sides of the traingle are : 6ft and 12ft.
in order to find the diagonal use the formula a^2 + b^2 = c^2
replacing values:
6ft^2 + 12ft^2 = c^2
36ft^2 + 144ft^2 = c^2
180ft^2 = c^2
c = squar root of 180ft^2
c = 13.42ft
therefore the diagonal should 13.42ft